OpenStudy (anonymous):

What is the surface area of a circle that has a diameter of 14 inches?

OpenStudy (anonymous):

If by surface area you mean area it will by pi(14/2)^2 otherwise it is going to be the circumference which equals 14pi

OpenStudy (anonymous):

ok these are the only options I have A. 49.86 in2 B. 153.86 in2 C. 196.86 in2 D. 616.86 in2

OpenStudy (anonymous):

since it is inches squared use the first formula for area and plug in 3.14 for pi

OpenStudy (anonymous):

The area of a circle is: \[\pi*r^2\] And the diameter is two radii. So the radius here is 7 inches. So \[\pi*(7)^2\]\[\pi*49\] That's about 153.94 (the exact answer using the exact value of pi in a calc) But here I'm assuming the directions are to substitute 3.14 for pi, since when you do that, you get 153.86, which is an answer choice. :)